OpenStudy (anonymous):

solve the equation by finding square roots 3x^2=21

OpenStudy (gorv):

x^2=21/3=7 x=7^(1/2)

OpenStudy (gorv):

@kimberly1097

OpenStudy (anonymous):

so its the square root of what

OpenStudy (gorv):

squareroot of 7

OpenStudy (gorv):

@kimberly1097

OpenStudy (anonymous):

but i got that answer wrong on my test

OpenStudy (gorv):

how u didi it????

OpenStudy (gorv):

show i will correct

OpenStudy (anonymous):

i plugged it in my calculator

OpenStudy (gorv):

solve it by urself

OpenStudy (anonymous):

i dont no how